# Magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances in hot and dense strange hadronic matter

### v1

@@ Abstract
This work comprises the calculations of magnetic moments of $\frac{1}{2}^{-}$ baryon resonances in the presence of hot and dense hadronic matter. Using the chiral $SU(3)$ quark mean field model, we have taken into account the impact on in-medium scalar meson fields to investigate the density effects on the in-medium baryon masses and their constituent quarks. In light of chiral constituent quark model $\chi$CQM, we have calculated the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances and scrutinized the explicit contributions coming from the valence quarks, sea quarks and the orbital angular momentum of sea quarks. The effective baryonic magnetic moments have been investigated for the varying values of isospin asymmetry and strangeness fraction.
I Introduction
@@
In quantum chromodynamics (QCD), quarks and gluons are treated as fundamental degrees of freedom which form bound states known as hadrons (baryons($qqq$) and mesons($q\bar{q}$)). These fundamental degrees of freedom undergo dynamical changes due to strong interactions in dense nuclear medium, which in turn cause in-medium modifications in the properties of hadrons. Confinement makes it challenging to understand QCD at lower energy scales. Scrutinizing the internal structure and the properties of light hadrons allow us to unravel the physics of strong interactions at low energies. Apart from heavy-ion collision experimental studies on chiral symmetry restoration at high temperature thirtyseven , and the de-confinement of hadrons to quark-gluon plasma (QGP) thirtyeight , it has become a primary interest for experiments such as PANDA at FAIR thirtythree ; thirtyfour , BESIII at BEPCII thirtysix ; thirtytwo , HADES at GSI thirtyfive etc., to explore the electromagnetic properties of hadrons in dense matter. Electromagnetic form factors, being key observable to investigate the electric and magnetic moment distributions, are of theoretical interest to explore other primary hadronic properties like magnetic moment and charged radii to probe deep inside of hadrons.
@@
Magnetic moment of hadrons, in free space, play a very important role in investigating their internal structure at a sub-nuclear level sixtyfour . Theoretically, the magnetic moments of octet as well as decuplet baryons have been extensively studied in free space one ; two ; three ; four ; five ; six ; seven ; eight ; nine . Isgur framed his model with one-gluon exchange, in order to account for internal structure of hadrons isgur1 ; isgur2 ; isgur3 ; isgur4 . The constituent quark model (CQM) propounded that the baryonic magnetic moments can be calculated as the summation of the individual magnetic moments of its constituent quarks ten ; eleven . It was Glozman who incorporated the Goldstone boson exchange (GBE) interaction in the low energy regime and the concept of one-gluon exchange was substituted by vector meson exchange interactions, which in fact led to better fits for baryon spectra glozman . It was not until the development of chiral constituent quark model ($\chi\mathrm{CQM}$) that the pseudo-scalar and scalar meson exchanges were considered which was later extended to the inclusion of vector meson exchange dai . Hence, $\chi\mathrm{CQM}$ was formed by extending the quark model and further incorporating the chiral symmetry and corresponding spontaneous symmetry breaking into it twelve . The $\chi\mathrm{CQM}$ model provides satisfactory explanation to the “Proton spin puzzle”, which was an outcome of the final data taking at EMC, and other hadron properties ten ; thirteen ; fourteen ; fifteen ; sixteen ; seventeen ; eighteen ; nineteen ; twenty ; twentyone ; twentytwo ; twentythree ; twentyfour ; twentyfive ; twentysix ; twentyseven . In previous studies, this model has been used for calculating the baryonic magnetic moments, by taking into account the sea quark polarizations and their orbital angular momentum contribution through the Cheng-Li mechanism twentyeight ; twentynine ; thirty . In the $\chi$CQM formalism, the magnetic moments of low-lying ${\frac{1}{2}}^{-}$ $N^{*}$ resonances thirtyone , ${\frac{1}{2}}^{-}$ octet baryon resonances thirtyone and low-lying ${\frac{1}{2}}^{-}$ $\Lambda$ resonances A_Torres have been calculated in earlier studies. Apart from $\chi$CQM, the magnetic moments of $N^{*}$ resonances have been computed by three quark Hamiltonian approach and the results were not consistent with lattice QCD values with very large variations hamiltonian_approach . As an extension to chiral perturbation theory, the Chiral unitary model was also used in ref. cum_hyodo to calculate the magnetic moments of $N^{*}(1535)$ resonances.
@@
As mentioned earlier, the study of magnetic moments of baryons and the medium dependent effects on the electromagnetic properties is an area of active research and stands relevant to some of the experiments. In order to study the properties of nuclear matter, various theoretical models like the MIT bag model thirtynine , linear sigma model Petro , non-linear sigma model fortyfive , Zimanyi and Moszkowski model seventyone , cloudy bag model fortythree , Nambu-Jona-Lasinio (NJL) model fortytwo , light-front quark model Schlumpf1993 , parity doublet model paritydoublet1 ; paritydoublet2 etc., have been developed. The cloudy bag model attained increased accuracy over MIT bag model in explaining QCD effects by incorporating the presence of pion clouds at the core of a baryon. With further development, models like the NJL model successfully explained low energy baryonic dynamics by considering chiral quarks as constituents of hadrons, through the inclusion of the interaction between the constituent quarks njlsuccess . Few years later, the quark meson coupling (QMC) model was constituted and it facilitated a deeper understanding of the in-medium baryon properties fortyfour . The QMC model, formulated in lines of cloudy bag and NJL model, has previously been deployed to study the impact of medium on the magnetic moments of octet baryons and has produced results which are in good agreement with experimental data seventyfour . The chiral $SU(3)$ quark mean field model (CQMF) fortyfour ; fortyfive ; fortysix ; fortyseven is an extended version of the quark mean field model fortyeight , which is based on the QMC model approach. In the CQMF model, the constituent quarks are confined into baryons by a confining potential based on mean field approximation using classical expectation values instead of quantum field operators fortynine . In the context of medium modifications of electromagnetic properties of hadrons, the studies in ref. kaon show the in-medium effects on $K^{+}$ meson properties in light-front quark meson coupling model (LF-QMC) in symmetric nuclear matter. With the rise in medium density, the kaon electromagnetic form factors decrease faster than in vacuum, while the kaon electromagnetic charge radii grow in size. By studying pion and kaon meson structures in NJL model the authors in ref. njlmeson have obtained results consistent with that in ref. kaon . For octet baryons in light of the QMC model effoctet , it has been shown that the in-medium electric form factors are dampened and the magnetic form factors are enhanced, and this finding is consistent with the in-medium variations of the bound proton electromagnetic form factors observed at Jefferson Lab jlab . The study of effective magnetic moments of baryons was also extended to calculate the low-lying charm, and the low-lying bottom baryons with the aim of understanding the roles of the heavy and light quarks in symmetric nuclear matter tsushima1 ; tsushima2 . While a comparative study on octet baryons with modified quark meson coupling model (MQMC) mqmc , suggests that the observed in-medium baryonic magnetic moments incur drastic variations compared to that in QMC framework. In CQMF formalism, the medium modifications of magnetic moments of octet fiftythree , decuplet fiftyfour and charmed Dutt:2024lui baryons have been studied in hot and dense hadronic matter. The CQMF model has recently been used to study valence quark distribution and transverse momentum dependent distributions of valence quarks of pions Puhan:2024xdq ; Kaur:2024wze .
@@
Presence of strangeness fraction ($f_{s}$) is a principal aspect to understand the departure of hadronic properties of baryons in strange hadronic matter comprising of hyperons from that in a non-strange medium. The in-medium variations of magnetic moments of octet and decuplet baryons in strange medium has shown stark deviation, as studied in fiftythree ; fiftyfour ; sixtytwo , for baryons of higher strangeness quantum number. In the present work, we have focused our study on the in-medium effects on the magnetic moments of the low-lying $N^{*}$ resonances and $\frac{1}{2}^{-}$ octet baryon resonances in CQMF model. It is exciting to study these resonances, as their detection connects hadron models to QCD properties in low and intermediate energy regimes. Photoproduction has been a key experimental tool involved in detecting the low-lying states of baryon resonances as shown in Refs. fifty ; fiftyone ; fiftytwo . The large decay widths of these excited baryon resonances, makes it challenging to detect them in experiments. In our current work, we have incorporated $\chi\mathrm{CQM}$ to calculate the properties of the baryon resonances as in thirtyone and have visualized them as a system of three valence quarks surrounded by the meson cloud. We have calculated the in-medium magnetic moments of the ${\frac{1}{2}}^{-}$ baryon resonances at finite temperature and density of hadronic medium through the medium modification of the corresponding constituent quark masses and the baryon masses. Within the CQMF framework, the effective masses of quarks ($m^{*}_{q}$) are calculated through the medium modified values of the scalar fields $\sigma$ and $\zeta$ and scalar iso-vector field $\delta$. The confined quarks inside the baryons, participate in the exchange interaction with these scalar meson fields as well as vector meson fields $\omega$, $\rho$ and $\phi$.
This manuscript has been organized as follows: In section II the CQMF model used to compute the in-medium quark masses and effective baryon masses, at finite temperature and density of the medium is presented. In section III we have calculated the contributions of valence quarks, sea quarks and orbital angular momentum of sea quarks towards the total magnetic moments of baryons, in light of $\chi\mathrm{CQM}$ model. Section IV is dedicated to a detailed discussion on the results obtained from numerical calculations. We devote section V to present the summary of the current work.
II Chiral Quark Mean Field Model
@@
In this section, we discuss briefly the framework, CQMF model, to calculate the in-medium constituent quark and baryon masses. As discussed earlier, in the CQMF model the quarks and mesons are treated as fundamental degrees of freedom, and the quarks are assumed to be confined inside baryons by an effective confining potential. The constituent quarks and mesons obtain their masses through the inclusion of properties such as spontaneous and explicit (for pseudoscalar mesons) symmetry breaking fortyfour . In the CQMF model, the effective Lagrangian density accounting for various interaction terms is given as fortyfour ; fortyfive ; fiftythree
${\cal L}_{{\rm eff}}\,=\,{\cal L}_{q0}\,+\,{\cal L}_{qm}\,+\,{\cal L}_{VV}\,+\,{\cal L}_{\Sigma\Sigma}\,+\,{\cal L}_{\chi SB}\,+\,{\cal L}_{c},$ (1)
@@
where $\,{\cal L}_{q0}=\bar{q}i\gamma^{\mu}\partial_{\mu}q\,$ is the kinetic term for the free part of massless quarks and ${\cal L}_{qm}\,$ is the quark-meson interaction term that remains invariant under chiral $SU(3)$ transformations. ${\cal L}_{qm}\,$ can be written as fortyfour ; fortysix ; fiftythree ; fiftyfour
$\mathcal{L}_{qm}$ $=g_{\mathrm{s}}\left(\bar{\Psi}_{L}M\Psi_{R}+\bar{\Psi}_{R}M^{\dagger}\Psi_{L}\right)-g_{v}\left(\bar{\Psi}_{L}\gamma^{\mu}l_{\mu}\Psi_{L}+\bar{\Psi}_{R}\gamma^{\mu}r_{\mu}\Psi_{R}\right)$ (2)
$=\frac{g_{\mathrm{s}}}{\sqrt{2}}\bar{\Psi}\left(\sum_{a=0}^{8}s_{a}\lambda_{a}+\mathrm{i}\gamma^{5}\sum_{a=0}^{8}p_{a}\lambda_{a}\right)\Psi-\frac{g_{v}}{2\sqrt{2}}\bar{\Psi}\left(\gamma^{\mu}\sum_{a=0}^{8}v_{\mu}^{a}\lambda_{a}-\gamma^{\mu}\gamma^{5}\sum_{a=0}^{8}a_{\mu}^{a}\lambda_{a}\right)\Psi.$
@@
In eq. 2, $\Psi=$ $\begin{pmatrix}u\\ d\\ s\end{pmatrix},$ and $g_{\mathrm{s}}$ and $g_{v}$ are the respective scalar and vector coupling constants. The relations between various quark-meson coupling constants are given as fortysix
@@
$\frac{g_{v}}{2\sqrt{2}}=g_{\rho}^{u}=-g_{\rho}^{d}=g_{\omega}^{u}=g_{\omega}^{d},~{}~{}~{}~{}~{}~{}~{}g_{\omega}^{s}=g_{\rho}^{s}=0.$ (4)
It is important to note that spin-0 scalar ($\Sigma$) and pseudoscalar ($\Pi$) meson nonets can be expressed as fortyfour ; fortysix
$M\left(M^{\dagger}\right)=\Sigma\pm i\Pi=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}\left(s^{a}\pm i\pi^{a}\right)\lambda^{a},$ (5)
where $s^{a}$ and $\pi^{a}$ represent the nonets of scalar and pseudoscalar mesons, respectively, while $\lambda^{a}$ are Gell-Mann matrices with $\lambda^{0}=\sqrt{\frac{2}{3}}I$. The spin-1 mesons can also be defined in a similar way as
$l_{\mu}\left(r_{\mu}\right)=\frac{1}{2}\left(V_{\mu}\pm A_{\mu}\right)=\frac{1}{2\sqrt{2}}\sum_{a=0}^{8}\left(v_{\mu}^{a}\pm a_{\mu}^{a}\right)\lambda^{a}.$ (6)
In the above expression, $v_{\mu}^{a}$ and $a_{\mu}^{a}$ are nonets of vector and pseudovector mesons, respectively. The physical states of scalar and vector meson nonets are given by
$\Sigma=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}s_{a}\lambda^{a}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\sigma+\delta^{0}\right)&\delta^{+}&\kappa^{*+}\\ \delta^{-}&\frac{1}{\sqrt{2}}\left(\sigma-\delta^{0}\right)&\kappa^{*0}\\ \kappa^{*-}&\bar{\kappa}^{*0}&\zeta\end{array}\right),$ (10)
@@
$V_{\mu}=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}v_{\mu}^{a}\lambda^{a}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\omega_{\mu}+\rho_{\mu}^{0}\right)&\rho_{\mu}^{+}&K_{\mu}^{*+}\\ \rho_{\mu}^{-}&\frac{1}{\sqrt{2}}\left(\omega_{\mu}-\rho_{\mu}^{0}\right)&K_{\mu}^{*0}\\ K_{\mu}^{*-}&\bar{K}_{\mu}^{*0}&\phi_{\mu}\end{array}\right),$ (14)
@@
The term, ${\cal L}_{VV}$ in eq. 1 represents the chiral-invariant vector meson self interaction terms. Within the framework of mean field approximation ${\cal L}_{VV}$ is given by fortyfour ; fiftythree ; fiftyfour
@@
${\cal L}_{\Sigma\Sigma}=$ $-\frac{1}{2}\,k_{0}\chi^{2}\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)+k_{1}\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)^{2}+k_{2}\left(\frac{\sigma^{4}}{2}+\frac{\delta^{4}}{2}+3\sigma^{2}\delta^{2}+\zeta^{4}\right)$
@@
The parameters $k_{0},k_{1},k_{2},k_{3}$ and $k_{4}$ appearing in the above Lagrangian density are calculated using masses of $\pi$ and $K$ as well as the average mass of $\eta$ and $\eta^{{}^{\prime}}$ mesons. The vacuum expectation values of scalar meson fields $\sigma$ and $\zeta$, i.e. $\sigma_{0}$ and $\zeta_{0}$, can be expressed in terms of pion and kaon leptonic decay constants $f_{\pi}$ and $f_{K}$, through the relations
@@
where $f_{\pi}=92.8$ MeV and $f_{K}=115$ MeV. The vacuum value of dilaton field $\chi_{0}=254.6$ MeV and the coupling constant $g_{4}=37.4$, obtained from fit reasonable effective nucleon mass fortyfour ; fiftythree ; fiftyfour . The last three terms of eq. 16 are introduced in the model to incorporate the trace anomaly property of QCD, which leads to the trace of energy momentum tensor being proportional to fourth power of dilaton field $\chi$ fortynine . The parameter $\xi$, originating from the logarithmic term used in eq. 16, is usually determined from QCD $\beta$-function at one loop level for three colors and flavors fortynine . The fifth term of the effective Lagrangian density of eq. 1 is the explicit chiral symmetry breaking term, ${\cal L}_{\chi SB}$, and is expressed as

### v2

# Magnetic moments of $\frac{1}{2}^{-}$ baryon resonances in strange hadronic matter at high densities

@@ Abstract
This work primarily focusses on determining the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances in the presence of hot and dense hadronic matter. In the chiral $SU(3)$ quark mean field model approach, we have essentially accounted for the effects on in-medium scalar meson fields to investigate the impact of high densities on the in-medium baryon masses and their constituent quarks. In light of chiral constituent quark model $\chi$CQM, we have calculated the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances and scrutinized the effects due to its internal constituents: the valence quarks, sea quarks and the orbital moment of sea quarks. Furthermore, we have investigated the effective baryonic magnetic moments for the finite magnitudes of isospin asymmetry and strangeness fraction.
[…]
xxxx, xxx
1 Introduction
@@
The theory of quantum chromodynamics (QCD), is based on the consideration of quarks and gluons as the fundamental degrees of freedom which form bound states known as hadrons (baryons($qqq$) and mesons($q\bar{q}$)). These fundamental degrees of freedom undergo dynamical changes due to strong interactions in dense nuclear medium, which in turn cause in-medium modifications in the properties of hadrons. Confinement makes it challenging to understand QCD at lower energy scales. Studying the internal structure and the properties of light hadrons allow us to unravel the physics of strong interactions at lower energy regimes. Apart from heavy-ion collision experimental studies on chiral symmetry restoration at high temperature thirtyseven , and the de-confinement of hadrons to quark-gluon plasma (QGP) thirtyeight , it has become a primary interest for experiments such as PANDA at FAIR thirtythree ; thirtyfour , BESIII at BEPCII thirtysix ; thirtytwo , HADES at GSI thirtyfive etc., to explore the electromagnetic properties of hadrons in dense matter. Electromagnetic form factors, being key observable to investigate the electric and magnetic moment distributions, are of theoretical interest to explore other primary hadronic properties like magnetic moment and charged radii to probe deep inside of hadrons.
@@
Magnetic moment of hadrons, in free space, play a very important role in investigating their internal structure at a sub-nuclear level sixtyfour . In theory, there has been extensive studies on the magnetic moments of octet and decuplet baryons in free space across a range of different models one ; two ; three ; four ; five ; six ; seven ; eight ; nine . Isgur framed his model with one-gluon exchange, in order to account for internal structure of hadrons isgur1 ; isgur2 ; isgur3 ; isgur4 . The constituent quark model (CQM) propounded that the baryonic magnetic moments are given as the summation of the individual magnetic moments of its constituent quarks ten ; eleven . It was Glozman who incorporated the Nambu Goldstone boson exchange (NGBE) interaction in the low energy regime and the concept of one-gluon exchange was substituted by vector meson exchange interactions, which in fact led to better fits for baryon spectra glozman . It was not until the development of chiral constituent quark model ($\chi\mathrm{CQM}$) that the pseudo-scalar and scalar meson exchanges were considered which was later extended to the inclusion of vector meson exchange dai . Hence, $\chi\mathrm{CQM}$ was formed by extending the quark model and further incorporating the chiral symmetry and corresponding spontaneous symmetry breaking into it twelve . The $\chi\mathrm{CQM}$ model provides satisfactory explanation to the “Proton spin puzzle”, which was an outcome of the final data taking at EMC, and other hadron properties ten ; thirteen ; fourteen ; fifteen ; sixteen ; seventeen ; eighteen ; nineteen ; twenty ; twentyone ; twentytwo ; twentythree ; twentyfour ; twentyfive ; twentysix ; twentyseven . In previous studies, this model has been used for calculating the baryonic magnetic moments, by taking into account the sea quark polarizations and their orbital angular momentum contribution through the Cheng-Li mechanism twentyeight ; twentynine ; thirty . In the $\chi$CQM formalism, the magnetic moments of low-lying ${\frac{1}{2}}^{-}$ $N^{*}$ resonances thirtyone , ${\frac{1}{2}}^{-}$ octet baryon resonances thirtyone and low-lying ${\frac{1}{2}}^{-}$ $\Lambda$ resonances A_Torres have been calculated in free space, in earlier studies. Apart from $\chi$CQM, the magnetic moments of $N^{*}$ resonances have been computed by three quark Hamiltonian approach and the results were not consistent with lattice QCD values with very large variations hamiltonian_approach . As an extension to chiral perturbation theory, the Chiral unitary model was also used in ref. cum_hyodo to calculate the magnetic moments of $N^{*}(1535)$ resonances.
@@
As mentioned earlier, the study of magnetic moments of baryons and the medium dependent effects on the electromagnetic properties is an area of active research and stands relevant to some of the experiments. For studying the properties of nuclear matter, various theoretical models like the MIT bag model thirtynine , linear sigma model Petro , non-linear sigma model fortyfive , Zimanyi and Moszkowski model seventyone , cloudy bag model fortythree , Nambu-Jona-Lasinio (NJL) model fortytwo , light-front quark model Schlumpf1993 , parity doublet model paritydoublet1 ; paritydoublet2 etc., have been developed. The cloudy bag model attained increased accuracy over MIT bag model in explaining QCD effects by incorporating the presence of pion clouds at the core of a baryon. With further development, models like the NJL model successfully explained low energy baryonic dynamics by considering chiral quarks as constituents of hadrons, through the inclusion of the interaction between the constituent quarks njlsuccess . Few years later, the quark meson coupling (QMC) model was constituted and it facilitated a deeper understanding of the in-medium baryon properties fortyfour . The QMC model, formulated in lines of cloudy bag and NJL model, has previously been deployed to study the impact of medium on the magnetic moments of octet baryons and has produced results which are in good agreement with experimental data seventyfour . The chiral $SU(3)$ quark mean field model (CQMF) fortyfour ; fortyfive ; fortysix ; fortyseven is an extended version of the quark mean field model fortyeight , which is based on the QMC model approach. In the CQMF model, the constituent quarks are confined into baryons by a confining potential based on mean field approximation using classical expectation values instead of quantum field operators fortynine . In the context of medium modifications of electromagnetic properties of hadrons, the studies in ref. kaon show the in-medium effects on $K^{+}$ meson properties in light-front quark meson coupling model (LF-QMC) in symmetric nuclear matter. With the rise in medium density, the kaon electromagnetic form factors decrease faster than in vacuum, while the kaon electromagnetic charge radii grow in size. By studying pion and kaon meson structures in NJL model the authors in ref. njlmeson have obtained results consistent with that in ref. kaon . For octet baryons in light of the QMC model effoctet , it has been shown that the in-medium electric form factors are dampened and the magnetic form factors are enhanced, and this finding is consistent with the in-medium variations of the bound proton electromagnetic form factors observed at Jefferson Lab jlab . There has been studies on the effective baryonic magnetic moments aimed to carry out calculation for the low-lying charm and the low-lying bottom baryons, to examine the roles of the heavy and light quarks in symmetric nuclear matter tsushima1 ; tsushima2 . While a comparative study on octet baryons with modified quark meson coupling model (MQMC) mqmc , suggests that the observed in-medium baryonic magnetic moments incur drastic variations compared to that in QMC framework. In CQMF formalism, the medium modifications of magnetic moments of octet fiftythree , decuplet fiftyfour and charmed Dutt:2024lui baryons have been studied in hot and dense hadronic matter. The CQMF model has recently been used to study valence quark distribution and transverse momentum dependent distributions of valence quarks of pions Puhan:2024xdq ; Kaur:2024wze .
@@
In the current work we calculate the magnetic moments of the low-lying $N^{*}$ resonances and $\frac{1}{2}^{-}$ octet baryon resonances in hot and dense hadronic matter. It is exciting to study these resonances, as their detection connects hadron models to QCD properties in low and intermediate energy regimes. Photoproduction has been a key experimental tool involved in detecting the low-lying states of baryon resonances as shown in refs. fifty ; fiftyone ; fiftytwo . The large decay widths of these excited baryon resonances, makes it challenging to detect them in experiments. While Ref. thirtyone emphasizes on magnetic moment calculations of negative parity octet baryons in free space, our study extends beyond free space in hot and dense hadronic medium. As we shall see later, to calculate the magnetic moments of constituent quarks the impact of relativistic quark confinement is taken into account in the calculations of present work fiftyfive . However, in Ref. thirtyone such confinement effects have not been considered and this causes a difference in the values of vacuum baryonic magnetic moments of this study as compared to our current calculations. Presence of strangeness fraction ($f_{s}$) is a principal aspect to understand the departure of hadronic properties of baryons in strange hadronic matter comprising of hyperons, from that in a non-strange medium. The in-medium variations of magnetic moments of octet and decuplet baryons in strange medium has shown stark deviation, as studied in refs. fiftythree ; fiftyfour ; sixtytwo , for baryons of higher strangeness quantum number. The effects of increasing isospin asymmetry has also been considered for the analysis of density dependent in-medium modifications of baryons. This manuscript has been organized as follows: Section 2 comprises of two subsections describing the theoretical models used in the current work. In section 2.1 the CQMF model used to compute the in-medium quark masses and effective baryon masses, at finite temperature and density of the medium is presented. In section 2.2 we have calculated the contributions of valence quarks, sea quarks and orbital angular momentum of sea quarks towards the total magnetic moments of baryons, in light of $\chi\mathrm{CQM}$ model. Section 3 is dedicated to a detailed discussion on the results obtained from numerical calculations. We devote section 4 to present the summary of the current work.
2 Theoretical Framework
Central to this work, we have used $\chi\mathrm{CQM}$ to calculate the properties of the baryon resonances as in Ref. thirtyone and have visualized them as a system of three valence quarks. We have calculated the in-medium magnetic moments of the ${\frac{1}{2}}^{-}$ baryon resonances at finite temperature and density of hadronic medium through the medium modification of the corresponding constituent quark masses and the baryon masses. Within the CQMF framework, the effective masses of quarks ($m^{*}_{q}$) are calculated through the medium modified values of the scalar fields $\sigma$ and $\zeta$ and scalar iso-vector field $\delta$. The confined quarks inside the baryons, participate in the exchange interaction with these scalar meson fields as well as vector meson fields $\omega$, $\rho$ and $\phi$.
2.1 Chiral Quark Mean Field Model
@@
In this section, the CQMF model is discussed briefly which serves as the basic framework to calculate the in-medium masses of constituent quark and baryon. As discussed earlier, in CQMF model the quarks and mesons serve as fundamental degrees of freedom, and the confinement of quarks by an effective confining potential results in the formation of baryons. The constituent quarks and the mesons obtain their masses through the inclusion of properties such as spontaneous and explicit (for pseudoscalar mesons) symmetry breaking fortyfour . In the CQMF model, the effective Lagrangian density accounting for various interaction terms is given as fortyfour ; fortyfive ; fiftythree
${\cal L}_{{\rm effective}}\,=\,{\cal L}_{q0}\,+\,{\cal L}_{qm}\,+\,{\cal L}_{VV}\,+\,{\cal L}_{\Sigma\Sigma}\,+\,{\cal L}_{\chi SB}\,+\,{\cal L}_{cp},$ (1)
@@
where $\,{\cal L}_{q0}=i\bar{q}\gamma^{\mu}\partial_{\mu}q\,$ represents the kinetic term for quarks having zero mass, while ${\cal L}_{qm}\,$ accounts for the interaction between quarks and mesons, which remains invariant when undergoing chiral $SU(3)$ transformations. ${\cal L}_{qm}\,$ is expressed as fortyfour ; fortysix ; fiftythree ; fiftyfour
$\mathcal{L}_{qm}$ $=g_{\mathrm{s}}\left(\bar{\Psi}_{R}{M_{n}}^{\dagger}\Psi_{L}+\bar{\Psi}_{L}{M_{n}}\Psi_{R}\right)-g_{v}\left(\bar{\Psi}_{L}\gamma^{\mu}L_{\mu}\Psi_{L}+\bar{\Psi}_{R}\gamma^{\mu}r_{\mu}\Psi_{R}\right)$ (2)
$=\frac{g_{\mathrm{s}}}{\sqrt{2}}\bar{\Psi}\left(\mathrm{i}\gamma^{5}\sum_{b=0}^{8}p_{b}\lambda_{b}+\sum_{b=0}^{8}s_{b}\lambda_{b}\right)\Psi-\frac{g_{v}}{2\sqrt{2}}\bar{\Psi}\left(\gamma^{\mu}\sum_{b=0}^{8}v_{\mu}^{b}\lambda_{b}-\gamma^{\mu}\gamma^{5}\sum_{b=0}^{8}a_{\mu}^{b}\lambda_{b}\right)\Psi.$
@@
In eq. 2, $g_{\mathrm{s}}$ and $g_{v}$ represent scalar and vector coupling constants, respectively and $\Psi=$ $\begin{pmatrix}u\\ d\\ s\end{pmatrix}$. The relationships between these coupling constants are given as fortysix
@@
$\frac{g_{v}}{2\sqrt{2}}=$ $g_{\rho}^{u}=-g_{\rho}^{d}=g_{\omega}^{u}=g_{\omega}^{d}=\frac{g_{\phi}^{s}}{\sqrt{2}},~{}~{}~{}~{}~{}~{}~{}g_{\omega}^{s}=g_{\rho}^{s}=g_{\phi}^{u}=g_{\phi}^{d}=0.$ (4)
It is important to note that spin-0 scalar, $\Sigma$ and pseudo-scalar meson nonets $\Pi$, are expressed as fortyfour ; fortysix
$M_{n}\left({M_{n}}^{\dagger}\right)=\Sigma\pm i\Pi=\frac{1}{\sqrt{2}}\sum_{b=0}^{8}\left(s^{b}\pm i\pi^{b}\right)\lambda^{b},$ (5)
where $s^{b}$ and $\pi^{b}$ represent scalar and pseudo-scalar meson nonents, respectively, while $\lambda^{b}$ represents Gell-Mann matrices with $\lambda^{0}=\sqrt{\frac{2}{3}}I$. The spin-1 mesons can also be defined by the relation
$L_{\mu}\left(r_{\mu}\right)=\frac{1}{2}\left(V_{\mu}\pm A_{\mu}\right)=\frac{1}{2\sqrt{2}}\sum_{b=0}^{8}\left(v_{\mu}^{b}\pm a_{\mu}^{b}\right)\lambda^{b}.$ (6)
In the above expression, $v_{\mu}^{b}$ and $a_{\mu}^{b}$ are vector and pseudovector meson nonents, respectively. The expressions representing scalar and vector meson nonets are
$\Sigma=\frac{1}{\sqrt{2}}\sum_{b=0}^{8}s_{b}\lambda^{b}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\delta^{0}+\sigma\right)&\delta^{+}&\kappa^{*+}\\ \delta^{-}&\frac{1}{\sqrt{2}}\left(\sigma-\delta^{0}\right)&\kappa^{*0}\\ \kappa^{*-}&\bar{\kappa}^{*0}&\zeta\end{array}\right),$ (10)
@@
$V_{\mu}=\frac{1}{\sqrt{2}}\sum_{b=0}^{8}v_{\mu}^{b}\lambda^{b}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\omega_{\mu}+\rho_{\mu}^{0}\right)&\rho_{\mu}^{+}&K_{\mu}^{*+}\\ \rho_{\mu}^{-}&\frac{1}{\sqrt{2}}\left(\omega_{\mu}-\rho_{\mu}^{0}\right)&K_{\mu}^{*0}\\ K_{\mu}^{*-}&\bar{K}_{\mu}^{*0}&\phi_{\mu}\end{array}\right),$ (14)
@@
The chiral invariant term ${\cal L}_{VV}$ in eq. 1, incorporates the mass term for the vector mesons as well as higher order self interactions and is written as fortyfour
${\cal L}_{VV}=\frac{1}{2}m_{V}^{2}\frac{\chi^{2}}{\chi_{0}^{2}}\operatorname{Tr}V_{\mu}V^{\mu}+2g_{4}\operatorname{Tr}\left(V_{\mu}V^{\mu}\right)^{2}.$ (15)
The expression for ${\cal L}_{VV}$ under mean field approximation comes out to be fortyfour ; fiftythree ; fiftyfour
@@
${\cal L}_{\Sigma\Sigma}=$ $-\frac{1}{2}\,k_{0}\chi^{2}\left(\delta^{2}+\zeta^{2}+\sigma^{2}\right)+k_{1}\left(\delta^{2}+\zeta^{2}+\sigma^{2}\right)^{2}+k_{2}\left(3\sigma^{2}\delta^{2}+\frac{\delta^{4}}{2}+\frac{\sigma^{4}}{2}+\zeta^{4}\right)$
@@
The parameters $k_{i=0,1,2,3,4}$ appearing in the above Lagrangian density are calculated using masses of $\pi$ and $K$ as well as the average of the $\eta$ and $\eta^{{}^{\prime}}$ meson masses. The expectation values of scalar meson fields $\sigma$ and $\zeta$, in free space i.e. $\sigma_{0}$ and $\zeta_{0}$, can be expressed with pion and kaon leptonic decay constants $f_{\pi}$ and $f_{K}$, with the expressions
@@
where $f_{\pi}=92.8$ MeV and $f_{K}=115$ MeV. In absence of nuclear matter, the dilaton field $\chi_{0}=254.6$ MeV and the coupling constant $g_{4}=37.4$, have been obtained from a reasonable fit of effective nucleon mass fortyfour ; fiftythree ; fiftyfour . The concluding three terms of eq. 17 incorporates the trace anomaly property of QCD, which leads to the trace of energy momentum tensor being proportional to the dilaton field $\chi$ raised to the power of 4 fortynine . The explicit chiral symmetry breaking term ${\cal L}_{\chi SB}$ in eq. 1 is expressed through the relation